A coil has a length of 3m having an electron density of 3.34 x 1024 e/m3 with a cross sectional area of 100 cm squared. How much current will pass through the conductor in 1hr?


1
Expert's answer
2021-10-05T17:04:01-0400

I. e I = vd​enawhere I = current vd​ =drift velocity e = electron n = no of electrons per unit volume a =area of cross section10=9×1028×10−4×vd​vd​=1.1×10−24